# Internal Consistency and Self-Feedback in Large Language Models: A Survey

### v2

@@ IV-F External Feedback
Sometimes, feedback from the model itself is not sufficient, and external feedback is needed. For example, in code generation, if there are hallucinations (bugs) in the code, it is difficult for even humans to accurately identify some bugs without executing the code with an external executor. Self-Debug [76] proposes using the execution results from an external executor as feedback. Besides using external tools, some works use other models as external feedback sources, such as a more powerful teacher model [77] or a peer model [73]. The commonly used RAG method, which can incorporates information retrieved from external sources as external feedback, is another example. However, as mentioned in Section I-E, using external information sources is not the focus of this paper as it does not pertain to Internal Consistency Mining. Nevertheless, many studies employing external signals in Self-Feedback are summarized in Section VII.
@@ V-A Reasoning Topologically
When answering a question, LLMs may employ different reasoning methods to solve the same problem. For instance, solving a system of linear equations can be done using either substitution elimination or addition-subtraction elimination. Similarly, LLMs may choose different reasoning paths, but not all reasoning paths lead to the correct answer. Therefore, finding reasoning paths that are consistent with the learned knowledge becomes a key issue, leading to a series of works focusing on optimizing reasoning paths. Fig. 8 summarizes the similarities and differences of these works.
@@
A survey [26] comprehensively examines various X-of-Thought methods, including Chain-of-Thought, Self-Consistency, Tree-of-Thought, and Graph-of-Thought. Here, we briefly introduce these methods. Readers can refer to [26] for more details. Input-Output (IO) is the most straightforward way to make a model reason. That is, ask a question and get an answer directly. However, due to the lack of intermediate steps, it often fails to solve complex problems. Consequently, the efficient and robust Chain-of-Thought (CoT) [10] was proposed, which requires the model to provide intermediate reasoning steps to avoid failures in solving difficult problems. However, this method’s limitation is that if there is an error in the reasoning, the final result will also be affected. Subsequently, Self-Consistency (SC) [2] was introduced, where a simple majority voting strategy can significantly improve the accuracy of the final answer. Yet, the majority voting strategy is inherently limited in its exploratory capabilities and can only address problems like QA, where there is a fixed answer set, allowing the selection of the most frequent answers from different responses. To address these limitations, Tree-of-Thought (ToT) [44] was proposed, which views reasoning as a path linking different thoughts, with each node having multiple successor nodes for thorough local exploration. This method also developed a general scoring system to let the language model evaluate the quality of nodes and decide whether to explore further. Graph-of-Thought (GoT) [45] further extended this line of work by providing aggregation among different thought nodes, enhancing the utilization of reasoning chains. Maieutic Prompting [78], similar to GoT, attempts to establish entailment relationships between thoughts through rules, then constructs a Max-SAT [79] problem to obtain the best choices.
Most X-of-Thought methods require sampling and aggregation of thoughts, often limited to queries with fixed label sets during aggregation. To solve this problem, several works have emerged. Multi-Perspective Self-Consistency (MPSC) [80] targets code generation tasks, evaluating each solution from multiple perspectives (solution, specification, and test case) to select the best one. Universal Self-Consistency (Universal SC) [81] uses LLMs instead of simple answer matching to choose the most selected response, enhancing the stability of the majority voting. Soft Self-Consistency (Soft SC) [75] proposes a more adaptive scoring function, calculating the joint probability of tokens in a response as the scoring function, thus extending the problem scope to soft labels.
Additionally, Quiet Self-Taught Reasoner (Quiet-STaR) [82] addresses the issue mentioned in Section II-B, where “although complex reasoning in responses is beneficial for solving intricate problems, they may disrupt model’s latent reasoning due to redundant reasoning text, thereby increasing response-level inconsistency.” Quiet-STaR samples rationales from model’s responses and wraps each rationale between special markers, that is, <|startofthought|> and <|endofthought|>, to assist next-token reasoning. These rationales are invisible to the user, making latent reasoning explicit and effectively reducing conflicts.
@@
Self-Update Strategy. For Self-Consistency prompting, the update uses majority voting result. For ToT prompting, the update method uses BFS and DFS strategies to search and select suitable thoughts as output. For GoT prompting, the update method is similar to ToT but includes more extensive search spaces, aggregating different thoughts.
@@ V-B Refining with Responses
Fine-tuning from the collected responses. This line of work involves “using self-generated data to fine-tune itself.” Specifically, they often use LLMs to produce multiple answers, select the better responses from them, and then use these better responses to fine-tune the model, enhancing its reasoning capabilities. For example, Self-Improve proposed by  [46] uses a majority voting strategy to obtain better outputs, collecting such data to fine-tune the model itself. Similarly, Tian et al. [88] propose a framework called Self-Improvement, which uses Monte Carlo Tree Search for data synthesis while generating fine-tuning datasets, improving model’s reasoning capabilities. This concept is not only effective in the reasoning domain but also finds applications in other fields like the web agent domain. Self-Improved Agents [89] improved performance by 31% using this method. In the preference optimization field, SRPO (Self-Improving Robust Preference Optimization) [90], and Self-Alignment [91] both utilize model-generated preferences to align with human preferences.
@@
Getting better response with NLI models. Besides fine-tuning methods, we also demonstrate rule-based optimization techniques using NLI. NLI is a classic task in traditional NLP that can determine the relationship between two statements as entailment / contradiction / neutral. With an NLI model, we can identify the relationships between multiple samples and find better responses. For instance, Agarwal et al. [93] use a pre-trained NLI model to identify and correct logically inconsistent statements generated by a pre-trained language model. They then convert the entailment and contradiction probabilities of the NLI into a Max-SAT problem [79], and use a constraint solver [94] to optimize and obtain more accurate and consistent predictions. Similarly, Mitchell et al. [47] propose ConCoRD (Consistency Correction through Relation Detection), which further utilizes a pre-trained NLI model to estimate the logical relationships between model predictions, constructs a factor graph representing the prediction probability distribution, and finds the maximum probability prediction result on the factor graph.
@@ V-C Multi-Agent Collaboration
Multi-Agent Debate [73] utilizes multiple peer models that engage in iterative debates, with a fixed number of rounds as the stopping condition. Their experiments show that debates with three or fewer rounds can generally lead to convergence among agents (i.e., LLMs consistently agreeing on the same answer). Xiong et al. [49] further propose the FORD (Formal Debate Framework), which introduces a Judge LLM to summarize the agents’ statements at the end, also using a fixed number of rounds as the stopping condition. They expand the scope of LLM debates by exploring the effects of debates among models with mismatched capabilities in various scenarios. REFINER [51] trains two models with different roles: a generator for intermediate reasoning steps and a critic for feedback, continuing the iterative dialogue until the correct answer is obtained or the critic has no further feedback. Notably, using the correct answer as a stopping condition has been criticized as unrealistic [96].
The Consensus Game proposed by Jacob et al. [97] deviates from the above frameworks by avoiding direct dialogue between LLMs. Instead, different LLMs participate in a game, based on the hypothesis that “asking a model for answer A to question Q (generative)” and “asking a model if A is the answer to Q (discriminative)” lack consistency [98]. To achieve consistency, they prompt the generator to produce both correct and incorrect answers, then use the discriminator to evaluate its own responses, aiming for the generator and discriminator to reach a consensus (Nash equilibrium). They select the best response based on the degree of consistency.
The significant drawback of this line of work is the high inference cost, as it often requires different LLM instances, potentially consuming multiple times the GPU memory and increasing the inference burden due to the extensive context generated by agents. Additionally, most models need a stopping condition to end the dialogue, and fixed round stopping is inflexible and can reduce performance. There is no current flexible and efficient stopping criterion. However, Multi-Agent systems remain a promising AI direction, and cost issues shouldn’t deter exploration. For instance, MACNet (Multi-Agent Collaboration Network) [50] uses dozens of agents and various network topologies to collaboratively solve problems. Despite the high costs, such exploration is beneficial for collaborative optimization.
@@ VI Task: Hallucination Alleviation
Reasoning elevation typically targets QA tasks, while hallucination alleviation is generally aimed at open-ended generation tasks such as story writing and code generation, emphasizing goals like fact enhancement, error reduction, and faithfulness enhancement. This section provides a survey of these directions. We have categorized four significant lines of work, as shown in the lower half of Table IV.
@@ VI-A Refining the Response Iteratively
This line of work is similar to Refining with Responses listed in Section V-B. The difference lies in the focus: Refining with Responses primarily targets QA tasks, where various methods often only need to consider intermediate reasoning steps and the correctness of the final answer. This can be achieved by sampling multiple responses and synthesizing or selecting the better ones. In contrast, hallucination alleviation primarily deals with open-ended tasks such as story generation and code generation. When sampling multiple responses, it is still necessary to meticulously check each response for errors. Therefore, to alleviate hallucinations, it is crucial to iteratively refine and polish a response, eliminating errors. Their comparison is shown in Table V.
TABLE V: Refining with Responses V.S. Refining the Response Iteratively
Section V-B Section VI-A
Name Refining with Responses Refining the Response Iteratively
Train. Half needed Few needed
Target Reasoning Elevation Hallucination Alleviation
Task QA: Math, NLI, etc. Open-Ended: Story, Code, etc.
Steps 1. Sample responses 2. Aggregate them 3. Refine 1. Generate one response 2. Refine it 3. Iterate
This line of work is relatively mature. The most famous include Self-Refine [8], Reflexion [52], and Self-Correct [9]. These three frameworks share the basic structure of having the LLM provide textual feedback, which is then used to update the response iteratively until a stopping criterion is met or the maximum iterations is reached, as shown in Algorithm 1.
@@
We can also formalize the process in Algorithm 1 as a conditional language model, as shown in Eq. 10.
$P(\bm{y}|\bm{x})=P(\bm{y}_{0}|\bm{x})\prod_{i=1}^{n}P(\bm{y}_{i}|\bm{y}_{i-1},\bm{x},f_{i-1})$ (10)
Here, $P(\bm{y}_{0}|\bm{x})$ is a common conditional language model, $n$ denotes the number of iterations, $f_{i}$ is the Feedback Signal, and $P(\bm{y}_{i}|\bm{y}_{i-1},\bm{x},f_{i-1})$ represents the refinement based on the previous output $\bm{y}_{i-1}$ and the generated feedback $f_{i-1}$.
@@ VI-B Mitigating Hallucination while Generating
Mündle et al. [21] utilize the phenomenon of Self-Contradiction to eliminate hallucinations [^13] Demo of Self-Contradiction: https://chatprotect.ai/. Specifically, it induces prompts to generate two contradictory sentences and then directs the LLM to resolve the contradictions, retaining the consistent information to generate a coherent response. Subsequent sentences follow a similar approach to produce a complete reply. Clearly, contradictory information is highly likely to be hallucinatory, thus effectively mitigating hallucinations. This method essentially extends Self-Consistency [2] into the domain of hallucination.
@@
In contrast to EVER, FAVA (FAct Vericaton with Augmentation) [54] adopts a more sophisticated approach. It fine-tunes the model to generate tokens that edit its own content, significantly enhancing editing efficiency. Their fine-tuning dataset includes examples like: “Messi is an <entity><delete >Argentine </delete ><mark >Brazilian </mark ></entity >soccer player.” Special tokens enclosed in angle brackets are also trained to be generated, effectively eliminating hallucinations through rendering. The major advantage of this method lies in granting the LLM maximum autonomy to make mistakes and subsequently correct them freely. Moreover, this approach bears resemblance to Quiet-STaR [82] mentioned in Section V-A, where both utilize special tokens to represent non-essential cognitive processes.
@@ VI-C Decoding Truthfully
The concepts of “Refining the response iteratively” and “Mitigating hallucination while generating” primarily aim to enhance response consistency. Conversely, Decoding Truthfully focuses predominantly on decoding consistency. In recent years, several studies have discovered that methods such as greedy decoding and sample decoding alone constrain LLMs from accurately expressing crucial information in natural language form within latent layers. Consequently, more complex and rational decoding strategies have been designed to elevate the reliability and accuracy of model’s responses.
@@
Unlike DoLa, SED [11] and DIVER  [57] focus on detecting and addressing discrepancies caused by differences in tokens at certain positions, termed Chaotic Points. Methods for detecting chaotic points include comparing the ratio of maximum to second-maximum token probabilities and ensuring the number of candidate tokens exceeds one, as shown in Eqs. 12 and  13, where $\delta_{r}$ is a probability threshold, $\gamma$ is a predefined coefficient, and $\mathcal{V}$ denotes the vocabulary. By assessing generated content against previously generated segments and potential tokens from chaotic points, scores such as information gain, weighted uncertainty, and weighted confidence help identify the most suitable token.
$\mathbb{I}_{1}=\mathbb{I}\left(P_{\text{gen}}\right)=\left(\frac{P_{\text{second }}}{\bm{p}_{\text{max }}}\geq\delta_{r}\right)$ (12)
$\tau$ $=\gamma\max_{w\in\mathcal{V}}P\left(w\mid\bm{y}^{0:t-1}\right)$ (13)
$\mathbb{I}_{2}$ =(|$=\left(\left|\left\{\bm{y}^{t}\in\mathcal{V}\mid P\left(\bm{y}^{t}\mid\bm{y}^{0:t-1}\right)\geq\tau\right\}\right|>1\right)$
Those methodologies primarily apply to closed-book generation tasks. For open-book generation tasks, current research focuses on leveraging external references to guide decoding. CAD [56] and ECAD [104] (named ECAD in this review) incorporate contextually relevant or irrelevant knowledge snippets into model inputs, intervening in the decoding process through contrastive decoding strategies to bridge the information gap between useful and non-useful information. To further exploit valuable information from references, FastMem [105] fine-tunes the final layer of LLMs to quickly memorize reference texts during inference, highlighting differences in decoding with and without references.
In summary, these studies identify internal inconsistency phenomena or instability in responses to guide LLMs’ probability distributions during decoding.
@@ VI-D Activating Truthfulness
Activating Truthfulness focuses on enhancing consistency in latent layers. Its core methods involve boosting attention heads and states that represent “truthfulness” within latent layers, or refining latent layers through fine-tuning, aiming to improve the model’s internal consistency.
@@
The exploration of latent truthfulness began with CCS (Contrast-Consistent Search) proposed by Burns et al. [106]. CCS investigates methods for mining knowledge embedded in latent layers by training a small classification head on Transformer latent layers. This method effectively activates model truthfulness, surpassing conventional inference methods.
@@
Inspired by CCS, scholars from Harvard University developed the Inference-Time Intervention (ITI) technique [5]. This technique involves two main steps: 1. Probe analysis: Using probe technology [^14] A probe is a small classifier where the input is latent states and the output is labels corresponding to a test task. In ITI, this test task is TruthfulQA [16], where a subset of test tasks trains the probe to identify attention heads associated with higher truthfulness (higher classification accuracy). to identify attention heads in the model related to truthfulness. 2. Inference-time intervention: Adjusting selected attention heads during the model’s answer generation process by increasing their weights. This guides the model towards more truthful reasoning paths. However, ITI has limitations in training probes using the last token’s latent layer state at the end of a question-answer pair. This approach lacks features and turns the problem into discerning truthfulness rather than generating truthful content (generative). Addressing these issues, TrFr [6] proposed the use of multi-dimensional orthogonal probes, effectively extracting features from both truthful and non-truthful texts to better identify effective attention heads. TruthX [7] explored a more efficient intervention strategy. It targets not only attention heads but also latent states in the forward feedback layer. Mapping these states separately using truthful and semantic encoders significantly reduces the impact on the language model’s overall performance while enhancing representations of truthfulness.
Besides improving truthfulness, what other characteristics do latent layers possess? Recently, more work has begun to explore various features within latent layers. This not only demystifies model black boxes but also offers potential methods to mitigate hallucination issues. For example, Wu et al. [107] discovered that some attention heads focus more on long-context retrieval capabilities (strong copy-paste needle abilities). In tests like Needle-in-a-Haystack, blocking these attention heads results in performance dropping from 94.7% to 63.6%. Can enhancing model retrieval heads reduce hallucinations in long contexts? This is a question worth exploring. For another example, teams from Anthropic also used dictionary learning to discover monosemanticity within LLMs [108], such as “Golden Gate Bridge.” Activating states corresponding to the “Golden Gate Bridge” in latent layers leads the model to consider itself as the bridge. OpenAI also proposed Sparse Autoencoder (SAE) [109] to automatically identify latent features within language models.
If feature exploration were more accessible, could we identifying “human-likeness” features that can prevent models from mistakenly perceiving themselves as human, thus reducing hallucinations related to cognitive errors? Exploring latent features in language models to subsequently reduce hallucinations from a white-box perspective is a promising direction. However, the main challenge lies in the inefficiency of probe technology, dictionary learning, SAE, and similar methods, which require training many classifiers to effectively extract latent features from models.
In addtion to implicit feature mining methods, Activating Truthfulness also includes truthfulness-oriented fine-tuning methods that essentially explore consistency within latent states. For instance, Tian et al. [110] proposed sampling multiple responses from an LLM and using Self-Evaluation to identify more factual responses. Collecting these QA pairs and retraining the original LLM aligns with the fine-tuning approach mentioned in Section V-B, though it is not the primary focus of this section.

### v3

@@ IV-F External Feedback
Sometimes, feedback from the model itself is not sufficient. For example, in code generation, if there are hallucinations (bugs) in the code, it is difficult for even humans to accurately identify some bugs without executing the code with an external executor. Self-Debug [49] proposes using the execution results from an external executor as feedback. Besides using external tools, some works use other models as external feedback sources, such as a more powerful teacher model [72] or a peer model [47]. The commonly used RAG method, which can incorporate information retrieved from external sources as external feedback, is another example utilizing external feedback.
@@ V-A Reasoning Topologically
When answering a question, LLMs may choose different reasoning paths, but not all reasoning paths lead to the correct answer. Therefore, finding reasoning paths that are consistent with the learned knowledge becomes a key issue, leading to a series of works focusing on optimizing reasoning paths. Fig. 8 summarizes the similarities and differences of these works.
@@
A survey [73] covers various X-of-Thought (XoT) methods. Input-Output (IO) is the simplest approach, asking a question and getting an answer directly, but often struggles with complex problems. To address this, Chain-of-Thought (CoT) [10] was introduced, adding intermediate reasoning steps, though errors in reasoning can affect results. Self-Consistency (SC) [2] improves accuracy via majority voting but is limited in exploratory power. Tree-of-Thought (ToT) [37] views reasoning as a path with multiple successor nodes for deeper exploration, while Graph-of-Thought (GoT) [38] aggregates reasoning chains across nodes. Similar to GoT, Maieutic Prompting [74] builds entailment relationships between thoughts, then constructs a Max-SAT [75] problem to obtain the best choices.
Most XoT methods require sampling and aggregation of thoughts, often limited to queries with fixed label sets during aggregation. To solve this problem, several works have emerged. Multi-Perspective Self-Consistency (MPSC) [76] targets code generation tasks, evaluating each solution from multiple perspectives (solution, specification, and test case) to select the best one. Universal Self-Consistency (Universal SC) [77] uses LLMs instead of simple answer matching to choose the most selected response, enhancing the stability of the majority voting. Soft Self-Consistency (Soft SC) [71] proposes a more adaptive scoring function, calculating the joint probability of tokens in a response as the scoring function, thus extending the problem scope to soft labels.
Additionally, Quiet Self-Taught Reasoner (Quiet-STaR) [39] addresses the issue mentioned in Section II-B, where “although complex reasoning in responses is beneficial for solving intricate problems, they may disrupt model’s latent reasoning due to redundant reasoning text, thereby increasing response-level inconsistency.” Quiet-STaR samples rationales from the model’s responses and wraps each rationale between special markers, that is, <|startofthought|> and <|endofthought|>, to assist next-token reasoning. These rationales are invisible to the user, making latent reasoning explicit and effectively reducing conflicts.
@@
Self-Update Strategy. For Self-Consistency prompting, the update uses a majority voting result. For ToT prompting, the update method uses BFS and DFS strategies to search and select suitable thoughts as output. For GoT prompting, the update method is similar to ToT but includes more extensive search spaces, aggregating different thoughts.
@@ V-B Refining with Responses
Fine-tuning from the collected responses. This line of work involves “using self-generated data to fine-tune itself.” Specifically, they often use LLMs to produce multiple answers, select the better responses from them, and then use these better responses to fine-tune the model, enhancing its reasoning capabilities. For example, Self-Improve [40] uses a majority voting strategy to obtain better outputs, collecting such data to fine-tune the model itself. Similarly, Tian et al. [83] propose a framework called Self-Improvement, which uses Monte Carlo Tree Search for data synthesis while generating fine-tuning datasets, improving model’s reasoning capabilities.
@@
Getting better response with NLI models. Besides fine-tuning methods, we also demonstrate rule-based optimization techniques using NLI [84, 41]. With an NLI model, we can identify the relationships between multiple samples and find better responses. For instance, Agarwal et al. [84] use a pre-trained NLI model to identify and correct logically inconsistent statements generated by a pre-trained language model. They then convert the entailment and contradiction probabilities of the NLI into a Max-SAT problem [75], and use a constraint solver [85] to optimize and obtain more accurate and consistent predictions.
@@ V-C Multi-Agent Collaboration
Debate Frameworks. Multi-Agent Debate [47] utilizes multiple peer models that engage in iterative debates, with a fixed number of rounds as the stopping condition. Their experiments show that debates with three or fewer rounds can generally lead to convergence among agents (i.e., LLMs consistently agreeing on the same answer). Xiong et al. [44] further propose the FORD (Formal Debate Framework), which introduces a Judge LLM to summarize the agents’ statements at the end, also using a fixed number of rounds as the stopping condition. They expand the scope of LLM debates by exploring the effects of debates among models with mismatched capabilities in various scenarios. REFINER [46] trains two models with different roles: a generator for intermediate reasoning steps and a critic for feedback, continuing the iterative dialogue until the correct answer is obtained or the critic has no further feedback. Notably, using the correct answer as a stopping condition has been criticized as unrealistic [87].
Game-Theoretic Approaches. The Consensus Game proposed by Jacob et al. [88] deviates from the above frameworks by avoiding direct dialogue between LLMs. Instead, different LLMs participate in a game, based on the hypothesis that “asking a model for answer A to question Q (generative)” and “asking a model if A is the answer to Q (discriminative)” lack consistency [89]. They prompt the generator to produce both correct and incorrect answers, then use the discriminator to evaluate its own responses, aiming for the generator and discriminator to reach a Nash equilibrium. They select the best response based on the degree of consistency.
The significant drawback of this line of work is the high inference cost, as it often requires different LLM instances, potentially consuming multiple times the GPU memory and increasing the inference burden due to the extensive context generated by agents. Additionally, most models need a stopping condition to end the dialogue, and fixed round stopping is inflexible and can reduce performance. There is no current flexible and efficient stopping criterion. However, Multi-Agent systems remain a promising AI direction, and cost issues shouldn’t deter exploration.
@@ VI Task: Hallucination Alleviation
Hallucination alleviation is aimed at open-ended generation tasks such as story writing and code generation, emphasizing goals like fact enhancement, error reduction, and faithfulness enhancement. We have categorized four significant lines of work, as shown in the lower half of Table IV.
@@ VI-A Refining the Response Iteratively
This line of work is similar to Refining with Responses (Section V-B) which primarily targets simple QA tasks. While Refining the Response Iteratively (Section VI-A) primarily deals with open-ended tasks such as story generation and code generation. Their comparison is shown in Fig. 9.
Figure 9: Refining with Responses (Left) V.S. Refining the Response Iter. (Right)
The most famous works include Self-Refine [8], Reflexion [48], and Self-Correct [9]. These three frameworks share the basic structure of having the LLM provide textual feedback, which is then used to update the response iteratively until a stopping criterion is met or the maximum iterations is reached, as shown in Algorithm 1.
@@ VI-B Mitigating Hallucination while Generating
Mündle et al. [18] utilize the phenomenon of Self-Contradiction to eliminate hallucinations [^9] Demo of Self-Contradiction: https://chatprotect.ai/. Specifically, it induces prompts to generate two contradictory sentences and then directs the LLM to resolve the contradictions, retaining the consistent information to generate a coherent sentence. Subsequent sentences follow a similar approach to produce a complete reply. Clearly, contradictory information is highly likely to be hallucinatory, thus effectively mitigating hallucinations. This method essentially extends Self-Consistency [2] into the domain of hallucination.
@@
In contrast to EVER, FAVA (FAct Vericaton with Augmentation) [51] adopts a more sophisticated approach. It fine-tunes the model to generate special tokens that edit its own content, enhancing editing efficiency [^10] Their fine-tuning dataset includes examples like: “Messi is an <entity><delete>Argentine </delete><mark>Brazilian </mark></entity >soccer player.” Special tokens enclosed in angle brackets are also trained to be generated, effectively eliminating hallucinations through rendering.. The major advantage of this method lies in granting the LLM maximum autonomy to make mistakes and subsequently correct them freely. Moreover, this approach bears resemblance to Quiet-STaR [39] mentioned in Section V-A, where both utilize special tokens to represent essential cognitive processes.
@@ VI-C Decoding Truthfully
Decoding Truthfully focuses predominantly on decoding consistency. In recent years, several studies have discovered that methods such as greedy decoding and sampling decoding constrain LLMs from accurately expressing crucial information in natural language. Consequently, more complex and rational decoding strategies have been designed to elevate the reliability and accuracy of model’s responses [94].
@@
Unlike DoLa, SED [11] and DIVER  [54] focus on detecting and addressing discrepancies caused by differences in tokens at certain positions, termed Chaotic Points. Methods for detecting chaotic points include comparing the ratio of maximum to second-maximum token probabilities or the number of candidate tokens exceeds one. Their indicator functions are shown in Eqs. 11 and  12, where $\delta_{r}$ is a probability threshold, $\gamma$ is a predefined coefficient, and $\mathcal{V}$ denotes the vocabulary. By assessing previously generated contents against potential tokens from chaotic points, scores such as information gain, weighted uncertainty, and weighted confidence help identify the most suitable token.
$\mathbb{I}_{1}\left(\frac{P_{\text{second }}}{\bm{p}_{\text{max }}}\geq\delta_{r}\right)$ (11)
$\mathbb{I}_{2}\left(\left|\left\{\bm{y}^{t}\mid P\left(\bm{y}^{t}\mid\bm{y}^{0:t-1}\right)\geq\gamma\max_{w\in\mathcal{V}}P\left(w\mid\bm{y}^{0:t-1}\right)\right\}\right|>1\right)$ (12)
Those methodologies primarily apply to closed-book generation tasks. For open-book generation tasks, current research focuses on leveraging external references to guide decoding. CAD [53] and ECAD [96] (named ECAD in this survey) incorporate contextually relevant or irrelevant knowledge snippets into model inputs, intervening in the decoding process through contrastive decoding strategies to bridge the information gap between useful and non-useful information.
@@ VI-D Activating Truthfulness
Activating Truthfulness focuses on enhancing consistency in latent layers. Its core methods involve boosting attention heads and states that represent “truthfulness” within latent layers, aiming to improve the model’s internal consistency.
@@
The exploration of latent truthfulness began with CCS (Contrast-Consistent Search) [97]. CCS investigates methods for mining knowledge embedded in latent layers by training a small classification head on Transformer latent layers. This method effectively activates model truthfulness, surpassing conventional inference methods.
@@
Inspired by CCS, Harvard scholars introduced the Inference-Time Intervention (ITI) technique [5]. ITI consists of two steps: 1) Probe analysis: Using probe technology [^11] A probe is a small classifier whose input is latent states and whose output is labels corresponding to a test task. to identify attention heads in the model related to truthfulness. 2) Inference-time intervention: The model’s answer generation process is adjusted by increasing the weights of selected attention heads, guiding the model toward more truthful reasoning. However, ITI has limitations in training probes using only the last token’s latent layer state at the end of a QA pair. TrFr [6] addressed this by using multi-dimensional orthogonal probes to extract features from both truthful and non-truthful texts, improving attention head identification. TruthX [7] explored a more efficient intervention strategy. It targets not only attention heads but also the feed-forward network layers. Mapping these states separately using truthful and semantic encoders significantly reduces the impact on the language model’s overall performance while enhancing representations of truthfulness.
White-Box Hallucination Alleviation. Mitigating hallucinations from a white-box perspective involves activating the internal authenticity of the model, which necessitates interpretability studies. For instance, a recent survey [98] reveals that attention heads in models can serve various functions. Building on these functional distinctions, we may discover better approaches to mitigate hallucinations. For example, Wu et al. [99] found that certain attention heads are more adept at long-context retrieval (strong “copy-paste” abilities). In tests such as Needle-in-a-Haystack, blocking these attention heads caused performance to drop from 94.7% to 63.6%. Can enhancing retrieval heads reduce hallucinations in long contexts? This is a question worth investigating.